Write an equation and solve the problem.
Find the sum of three consecutive odd integers if the sum of the first two integers is equal to 24 less than four times the third integer.

Answer by oberobic(2304)

Three consecutive odd integers can be defined as:
x
x+2
x+4
.
x + x+2 = sum of first two numbers = 2x+2
.
2x+2 = 4(x+4) -24
.
2x +2 = 4x + 16 -24
.
-2x = -10
.
x = 5
so
x+2 = 7
x+4 = 9
.
Check the answer.
.
4*9 = 36
5+7 = 12
36-24=12
OK
.
The three numbers are 5, 7, and 9.